OpenStudy (anonymous):

a rocket is shot vertically into the air so that its height above ground after t seconds is h(t)=180t + 30t^2 - 40t^3 feet. At what time is the velocity of the rocket at its maximum? What is its maximum distance above ground?

OpenStudy (anonymous):

Dh(t) = 180 + 60t -120t^2 = 3 + t -2t^2 = (3-2t)(1+t) = 0 for maxt -> t =2/3 and f (2/3) ¨= maxs = 405/2 feet.

OpenStudy (anonymous):

i got 2/3 too! but the answer key says 1/4 so i was thrown off

OpenStudy (anonymous):

Sorry, typo, 3/2

OpenStudy (anonymous):

same here...i still go the same answer as you.